Evaluate each expression
sin 5 pie over 6
tan 5 pie over 4

Answer by stanbon(75887)

Evaluate each expression
sin 5 pie over 6
----
sin(5pi/6) = sin(pi/6) = 1/2
-------------------------------
tan 5 pie over 4
-----
tan(5pi/4) = tan(pi/4) = 1
=============================
Cheers,
Stan H.